OpenStudy (anonymous):

Which of the following is equal to C(10, 3)? a. C(3, 10) b. C(10, 7) c. P(10, 3)

OpenStudy (anonymous):

the relation between combination and permutation is ⁿPr = (r!) · ⁿCr i hope you could do so

OpenStudy (lgbasallote):

\[C(10,3) \implies \frac{10!}{3! (10-3)!}\] does that help?

OpenStudy (anonymous):

thats P(10,3)

OpenStudy (anonymous):

remember nCr = nC(n-r) what is n-r?

OpenStudy (anonymous):

See how the graph is symmetrical about x = 5? this is where x! = (10 - x)!

OpenStudy (anonymous):

omg Mac :P nice

OpenStudy (anonymous):

@apple_pi

OpenStudy (anonymous):

Mac ftw!